OpenStudy (anonymous):

solve the equation 6n+11=-7+3n for n

OpenStudy (anonymous):

help!!!

OpenStudy (anonymous):

n=-6

OpenStudy (anonymous):

6n-3n=-7-11 3n=-18 n=-6